# Efficient Offline Reinforcement Learning: The Critic is Critical

### v1

@@ Abstract
Recent work has demonstrated both benefits and limitations from using supervised approaches (without temporal-difference learning) for offline reinforcement learning. While off-policy reinforcement learning provides a promising approach for improving performance beyond supervised approaches, we observe that training is often inefficient and unstable due to temporal difference bootstrapping. In this paper we propose a best-of-both approach by first learning the behavior policy and critic with supervised learning, before improving with off-policy reinforcement learning. Specifically, we demonstrate improved efficiency by pre-training with a supervised Monte-Carlo value-error, making use of commonly neglected downstream information from the provided offline trajectories. We find that we are able to more than halve the training time of the considered offline algorithms on standard benchmarks, and surprisingly also achieve greater stability. We further build on the importance of having consistent policy and value functions to propose novel hybrid algorithms, TD3+BC+CQL and EDAC+BC, that regularize both the actor and the critic towards the behavior policy. This helps to more reliably improve on the behavior policy when learning from limited human demonstrations. Code is available at: https://github.com/AdamJelley/EfficientOfflineRL
@@ 1 Introduction
Recent work has highlighted the effectiveness of supervised learning approaches (without temporal difference learning) for offline reinforcement learning (Emmons et al.,, 2022; Chen et al.,, 2021; Brandfonbrener et al.,, 2021; Peng et al.,, 2019). Other work has analysed in detail the limitations of these supervised approaches and when off-policy reinforcement learning techniques should be favoured (Kumar et al., 2022b, ; Brandfonbrener et al.,, 2022; Paster et al.,, 2022). Given these seemingly opposing approaches, it is natural to ask: can we get the best of both supervised learning and temporal difference learning for offline reinforcement learning? Specifically, can we provide an approach that provides the training efficiency and stability of supervised learning, while still gaining the performance benefits of multi-step temporal difference learning?
@@
In this work we investigate such an approach, by supervised pre-training off-policy reinforcement algorithms to obtain approximate behavior policy and values before attempting improvement. Our core contribution is to demonstrate that pre-training the critic in addition to the policy is crucial to achieve efficient and stable improvement beyond the performance of the behavior policy.
As a motivational example, we can consider offline tabular $Q$-learning on the simple 4 state MDP illustrated below, where we are provided with a single offline trajectory from an unknown policy.
@@
Figure 1: A motivational tabular MDP. In offline reinforcement learning, we are provided with a dataset of trajectories. In this paper we utilize information from the entire trajectory to help initialize a self-consistent critic for off-policy reinforcement learning, which eliminates much of the initial inefficiency and instability associated with bootstrapping in temporal difference losses.
@@
Tabular $Q$-learning initializes all $Q$-values to zero (or equivalently $Q$-networks are randomly initialized such that all initial $Q$-values are close to zero for deep $Q$-learning), and then performs temporal difference updates using the following update rule (Sutton et al.,, 2018):
@@
0 0 0 0 0 0 or +0.5111The MC initialization value for $Q(0,\rightarrow)$ depends on the choice of first-visit or every-visit Monte-Carlo (Sutton et al.,, 2018). In this example it doesn’t matter which is used after initialization. 0 +1 +3
@@
This minimal example demonstrates that the benefit of pre-training values on offline data arises even in the absence of function approximation, near-optimal behavior policies, or complex MDPs. However, the inefficiency of boostrapping from randomly initialized temporal difference (TD) targets still exists in the case of uninformed neural network initializations, and in sparse-reward environments where rewards received at the end of long trajectories may take many TD updates to propagate.
In this paper, we provide similar approaches for pre-training a critic or value network, which can be combined with most existing off-policy reinforcement learning algorithms, leading to improved training efficiency and stability. We first consider return-maximising reinforcement learning, and apply our approach to a minimal offline reinforcement learning baseline, TD3+BC (Fujimoto and Gu,, 2021). We then generalize our procedure for entropy-regularized reinforcement learning, and apply our approach to a state-of-the-art offline baseline, EDAC (An et al.,, 2021). Finally, we build on our insight into the importance of having consistent policy and value functions by introducing novel hybrid algorithms, TD3+BC+CQL and EDAC+BC, that regularize both the actor and the critic towards the behavior policy and values to enable smooth performance improvement over the pre-training performance on the challenging Adroit environments (Rajeswaran et al.,, 2018).
2 Preliminaries and Related Work
Perhaps the most straightforward approach for learning a policy from the offline data is behavior cloning (Pomerleau,, 1991). Since the training is supervised, convergence is relatively stable and efficient, but the learned policy can at best match the performance of the behavior policy since behavior cloning does not utilize reward information, and online performance may be brittle due to accumulating errors taking the agent out-of-distribution (OOD) of known states (Ross et al.,, 2011).
What if we want to improve on the behavior policy? Behaviour cloning variants such as BC-$k\%$ (Levine et al.,, 2020) and Advantage-Weighted Regression (Peng et al.,, 2019; Peters and Schaal,, 2007) utilize reward information to selectively clone the behavior data. More recently, conditioning the policy on desired returns or goals (Srivastava et al.,, 2021; Ma et al.,, 2022) has seen some success with transformer based approaches (Chen et al.,, 2021; Janner et al.,, 2021; Carroll et al.,, 2022). While these behavior cloning variants are sometimes able to achieve generalization to greater returns at test-time than observed in the dataset (Brandfonbrener et al.,, 2022; Kumar et al., 2022b, ), in general their performance is more limited than true reinforcement learning approaches which can more effectively use mechanisms such as trajectory stitching (Paster et al.,, 2022; Yang et al.,, 2022).
A more promising approach for improving on the behavior policy is to use off-policy reinforcement learning, usually in the form of actor-critic algorithms (Lillicrap et al.,, 2019; Silver et al.,, 2014). However, naively taking the offline dataset as the replay buffer for an off-policy algorithm usually leads to policy collapse (Fujimoto et al.,, 2019). This occurs because as the policy (actor) and values (critic) are optimized for maximal return, they necessarily go out-of-distribution of the data (Chen and Jiang,, 2019). Since there are no additional interactions to provide correcting feedback on these actions and values as in the online case, growing extrapolation errors cause erroneous values and actions that lead to performance equivalent to random. Most modifications of off-policy reinforcement learning algorithms for the offline setting involve regularization of either the actions or the values towards the provided dataset to prevent this out-of-distribution extrapolation (Fu et al.,, 2022). TD3+BC (Fujimoto and Gu,, 2021) modifies TD3 (Fujimoto et al.,, 2018) by introducing a behavior cloning term to regularize the policy towards the behavior policy. Alternatively, CQL (Kumar et al.,, 2020) modifies Q-learning to regularize the values for out-of-distribution actions to prevent positive extrapolation error. However, since regularization towards the behavior policy or values limits performance improvement (Moskovitz et al.,, 2022), recent approaches instead aim to capture out-of-distribution uncertainty (Wu et al.,, 2021). SAC-N and EDAC (An et al.,, 2021) use the minimum of an ensemble of critics to obtain value estimates that minimise positive extrapolation error (with EDAC introducing additional diversification loss over SAC-N to reduce the required ensemble size), such that policy optimization is less likely to lead to policy collapse.
While these off-policy RL approaches can lead to better performance than modified imitation approaches, their convergence can be inefficient due to the bootstrapping in the Bellman update used for learning values. In this work we attempt to address these difficulties by pre-training with supervised objectives, utilizing Monte-Carlo (MC) return estimates. The use of MC values has a long history in reinforcement learning, for which additional related work is discussed in Appendix A.
3 Theoretical Motivation
@@
In this section we provide a more formal analysis of the intuition provided for the minimal tabular $Q$-learning example in the introduction. We aim to investigate the functional form with which the $Q$-function initialization can affect the convergence to optimal $Q$-values. We consider the case of approximate fitted $Q$-iteration (Riedmiller,, 2005), in which:
$\hat{Q}_{k+1}\leftarrow\arg\min_{\hat{Q}}||\hat{Q}-\hat{T}\hat{Q}_{k}||_{\infty}$ (3)
@@
where we have defined $\hat{Q}_{k}$ as the estimate of the $Q(s,a)$ function after $k$ iterations of Equation 3, and $\hat{T}$ as the approximate Bellman operator $\hat{T}Q=\hat{r}+\gamma\hat{P}\max_{a}Q$, where $\hat{r}(s,a)$ and $\hat{P}(s^{\prime}|s,a)$ are the sampled reward and environment transition probabilities respectively. The exact equivalents are denoted without the circumflex. We consider minimizing the infinite norm for each iteration since there is no convergence guarantee for the squared norm commonly used for empirical loss functions (Agarwal et al.,, 2022). In general, the error due to each $Q$-iteration (between successive iterations of the $\hat{Q}$ function) can be written as the sum of an ‘approximation’ error, and a ‘sampling’ error:
$||\hat{Q}_{k}-T\hat{Q}_{k-1}||_{\infty}$ $=||\hat{Q}_{k}-\hat{T}\hat{Q}_{k-1}+\hat{T}\hat{Q}_{k-1}-T\hat{Q}_{k-1}||_{\infty}$ (4)
$\leq\underbrace{||\hat{Q}_{k}-\hat{T}\hat{Q}_{k-1}||_{\infty}}_{\text{approximation}}+\underbrace{||\hat{T}\hat{Q}_{k-1}-T\hat{Q}_{k-1}||_{\infty}}_{\text{sampling}}$ (5)
Since the offline data is provided for offline reinforcement learning, the samples from which the Bellman operator can be approximated is fixed. We additionally assume here that the entire dataset is sampled uniformly to convergence in each iteration, such that the sampling error is constant and unaffected by $Q$-function initialization. We therefore focus on the approximation error, and assume this is finite at each iteration, $||\hat{Q}_{k+1}-T\hat{Q}_{k}||_{\infty}\leq||\hat{Q}_{k+1}-\hat{T}\hat{Q}_{k}||_{\infty}\leq\epsilon_{k}$. Since we are interested in the effect of $Q$-function initialization on final optimality, we write (Agarwal et al.,, 2022):
$||\hat{Q}_{k}-Q^{*}||_{\infty}$ $=||\hat{Q}_{k}-T\hat{Q}_{k-1}+T\hat{Q}_{k-1}-Q^{*}||_{\infty}$ (6)
$=||(\hat{Q}_{k}-T\hat{Q}_{k-1})+(T\hat{Q}_{k-1}-T{Q^{*}})||_{\infty}$ (7)
$\leq||\hat{Q}_{k}-T\hat{Q}_{k-1}||_{\infty}+||T\hat{Q}_{k-1}-T{Q^{*}}||_{\infty}$ (8)
$\leq\epsilon_{k-1}+||T\hat{Q}_{k-1}-T{Q^{*}}||_{\infty}$ (9)
$\leq\epsilon_{k-1}+\gamma||\hat{Q}_{k-1}-Q^{*}||_{\infty}$ (10)
$\leq\sum_{i=0}^{k-1}\gamma^{i}\epsilon_{k-i-1}+\gamma^{k}||\hat{Q}_{0}-Q^{*}||_{\infty}$ (11)
@@
This provides a measure of the (inifite-norm) Bellman error after $k$ iterations of fitted $Q$-iteration. We assume that for final convergence to the optimal policy, we would like to achieve an error over the offline dataset of order $\delta$. Therefore for two different intializations of $Q$, $Q_{0}$ and $Q^{\prime}_{0}$, we assume that it takes $k$ and $k^{\prime}$ iterations respectively to reach this error level, and make the bound tight, so we have:
$\delta=\sum_{i=0}^{k-1}\gamma^{i}\epsilon_{k-i-1}+\gamma^{k}||\hat{Q}_{0}-Q^{*}||_{\infty}$ (12)
$\delta=\sum_{i=0}^{k^{\prime}-1}\gamma^{i}\epsilon^{\prime}_{k^{\prime}-i-1}+\gamma^{k^{\prime}}||\hat{Q^{\prime}}_{0}-Q^{*}||_{\infty}$ (13)
@@
$\sum_{i=k^{\prime}}^{k-1}\gamma^{i}\epsilon$ $=\gamma^{k^{\prime}}||\hat{Q^{\prime}}_{0}-Q^{*}||_{\infty}-\gamma^{k}||\hat{Q}_{0}-Q^{*}||_{\infty}$ (14)
$\frac{\epsilon(\gamma^{k^{\prime}}-\gamma^{k})}{1-\gamma}$ $=\gamma^{k^{\prime}}||\hat{Q^{\prime}}_{0}-Q^{*}||_{\infty}-\gamma^{k}||\hat{Q}_{0}-Q^{*}||_{\infty}$ (15)
$k-k^{\prime}$ $=\frac{1}{\ln\gamma}\ln\left(\frac{\frac{\epsilon}{1-\gamma}-||\hat{Q^{\prime}}_{0}-Q^{*}||_{\infty}}{\frac{\epsilon}{1-\gamma}-||\hat{Q}_{0}-Q^{*}||_{\infty}}\right)$ (16)
We see that $\hat{Q}^{\prime}_{0}\rightarrow\hat{Q}_{0}\implies k^{\prime}\rightarrow k$, i.e. a given initialization requires a given number of fitted $Q$-iterations to converge, as expected. We also see that as $\hat{Q}^{\prime}_{0}$ becomes closer to $Q^{*}$ (in terms of the infinite norm over all states and actions), the numerator grows with respect to the denominator and so $k-k^{\prime}$ increases logarithmically, demonstrating that the number of fitted $Q$-iterations $k^{\prime}$ decreases as the $Q^{\prime}_{0}$ initialization improves.
In this work, we parameterize the $Q$ function with a network and assume this can be initialized to the approximate behavior value function $Q_{B}$ by pre-training the network with a supervised Monte-Carlo value error, so $\hat{Q}^{\prime}_{0}=\hat{Q}_{B}$. If we assume that the behavior policy $\pi_{B}$ that generated the offline data achieves returns better than random (zero), we expect that $||\hat{Q}_{B}-Q^{*}||_{\infty}<||Q^{*}||_{\infty}$. If this holds, then Equation 16 demonstrates that the training efficiency (number of $Q$-iterations to convergence) should improve by initializing $\hat{Q}^{\prime}_{0}=\hat{Q}_{B}$ compared to $\hat{Q}_{0}=0$. However, there are many ways in which this theory could be deviated from in practice, such as the use of generalized value iteration (rather than fitted $Q$-iteration) used by modern deep actor-critic methods, the use of a squared- rather than infinite-norm loss function, the strong assumption that the error in each fitted $Q$-iteration is constant for any initialization, and the fact that variance in Monte-Carlo value targets may be such that $||\hat{Q}_{B}-Q^{*}||_{\infty}\not<||Q^{*}||_{\infty}$ (especially given the infinite norm). Despite these limitations, this theoretical analysis provides clear motivation for our practical implementation outlined below.

### v2

# Efficient Offline Reinforcement Learning: First Imitate, then Improve

@@ Abstract
Supervised imitation-based approaches are often favored over off-policy reinforcement learning approaches for learning policies offline, since their straightforward optimization objective makes them computationally efficient and stable to train. However, their performance is fundamentally limited by the behavior policy that collected the dataset. Off-policy reinforcement learning provides a promising approach for improving on the behavior policy, but training is often computationally inefficient and unstable due to temporal-difference bootstrapping. In this paper, we propose a best-of-both approach by pre-training with supervised learning before improving performance with off-policy reinforcement learning. Specifically, we demonstrate improved efficiency by pre-training an actor with behavior cloning and a critic with a supervised Monte-Carlo value error. We find that we are able to substantially improve the training time of popular off-policy algorithms on standard benchmarks, and also achieve greater stability. Code is available at: https://github.com/AdamJelley/EfficientOfflineRL.
@@ 1 Introduction
Supervised imitation-based approaches (emmons_rvs_2022; chen_decision_2021; peng_advantage-weighted_2019) are becoming increasingly popular for learning policies offline due to their straightforward optimization objective and ease of use with large foundation models (kim_openvla_2024; zhao_learning_2023; brohan_rt-2_2023; brohan_rt-1_2022). However, recent work has highlighted the limitations of these supervised approaches and demonstrated that off-policy reinforcement learning (RL) leads to improved performance when the data is sub-optimal or noisy, and can be more reliable for stochastic environments and long-horizon tasks (kumar_when_2022; brandfonbrener_when_2022; paster_you_2022). Given these seemingly opposing approaches, it is natural to ask: can we get the best of both supervised learning and temporal-difference learning for offline reinforcement learning? Specifically, can we provide an approach that provides the training efficiency and stability of supervised learning, while still gaining the performance benefits of multi-step temporal difference learning?
@@
In this paper, we investigate such a supervised pre-training approach to obtain approximate behavior policies and value functions before attempting improvement with off-policy RL. While a naive approach might simply be to pre-train the policy with behavior cloning, this generally does not provide much benefit (as we will show), since it does not provide a value function, which off-policy RL algorithms require to improve the policy. Our core contribution is to demonstrate that pre-training a critic is essential to achieve efficient and stable improvement beyond the behavior policy performance. Our approach can be straightforwardly combined with existing off-policy reinforcement learning algorithms with minimal modification, to improve their training efficiency and stability.
The remainder of the paper is structured as follows. We begin by reviewing the relevant background for offline reinforcement learning. We then present a motivational example to provide intuition for why supervised pre-training can improve subsequent off-policy RL, before building on this example to provide a theoretical analysis of the expected efficiency improvement from pre-training. We generalize our procedure for entropy-regularized reinforcement learning, and provide algorithms for applying our approach in practice to modern offline RL algorithms. Finally, we demonstrate that the hypothesized efficiency gains hold in practice on standard offline RL benchmarks.
2 Preliminaries and Related Work
Perhaps the most straightforward approach for learning a policy from offline data is behavior cloning (BC) (pomerleau_efficient_1991). Since the training is supervised, convergence is relatively stable and efficient, but the learned policy can at best match the performance of the behavior policy, since behavior cloning does not utilize reward information, and online performance may be brittle due to accumulating errors taking the agent out-of-distribution (OOD) of known states (ross_reduction_2011).
What if we want to improve on the behavior policy? Behavior cloning variants such as BC-$k\%$ (levine_offline_2020) and Advantage-Weighted Regression (peng_advantage-weighted_2019; peters_reinforcement_2007) utilize reward information to preferentially clone the behavior data. More recently, conditioning the policy on desired returns or goals (srivastava_training_2021; ma_how_2022) has seen some success with transformer based approaches (chen_decision_2021; janner_offline_2021; carroll_unimask_2022). While these behavior cloning variants are sometimes able to achieve generalization to greater returns at test-time than observed in the dataset (brandfonbrener_when_2022; kumar_when_2022), in general their performance is more limited than true reinforcement learning approaches which can more effectively use mechanisms such as trajectory stitching (paster_you_2022; yang_dichotomy_2022).
A more promising approach to improve on the behavior policy is to use off-policy reinforcement learning, usually in the form of actor-critic algorithms for continuous environments (lillicrap_continuous_2019; silver_deterministic_2014). However, naively taking the offline dataset as the replay buffer for an off-policy algorithm often leads to performance no better than random (fujimoto_off-policy_2019). This occurs because as the actor and critic are optimized for maximal return, they inevitably go out-of-distribution of the data (chen_information-theoretic_2019). Since there are no additional interactions to provide correcting feedback as in the online case, growing extrapolation errors cause erroneous values, amplified by temporal-difference bootstrapping, which can lead to policy collapse (luo_finetuning_2023).
Most modifications of off-policy reinforcement learning algorithms for the offline setting involve regularization of either the actions or the values towards the provided dataset to prevent this out-of-distribution extrapolation (fu_closer_2022). TD3+BC (fujimoto_minimalist_2021) modifies TD3 (fujimoto_addressing_2018) by introducing a behavior cloning term to regularize the policy towards the behavior policy. Alternatively, CQL (kumar_conservative_2020) modifies Q-learning to reduce the values for out-of-distribution actions to prevent positive extrapolation error. However, since regularization towards the behavior policy or values limits performance improvement (moskovitz_tactical_2022), recent approaches instead aim to capture out-of-distribution uncertainty (wu_uncertainty_2021). SAC-N and EDAC (an_uncertainty-based_2021) use the minimum of an ensemble of critics to obtain value estimates that minimize positive extrapolation error (with EDAC introducing an additional diversification loss over SAC-N to reduce the required ensemble size), so optimization is less likely to cause policy collapse.
While these off-policy RL approaches can lead to better performance than modified imitation approaches, their convergence can be inefficient due to the bootstrapping in the Bellman update used for learning values. It is well established that the Bellman error can be a poor proxy for the real value error, particularly when used for incomplete, off-policy datasets (fujimoto_why_2022; patterson_generalized_2022; schulman_equivalence_2018), which causes difficulties with using the Bellman error as the sole objective for training value functions offline. In this work we attempt to address these difficulties by pre-training with a supervised objective, using Monte-Carlo (MC) return estimates.
Monte-Carlo (MC) return estimates have a long history in online reinforcement learning for integrating information over longer horizons (sutton_reinforcement_2018; he_learning_2016; ostrovski_count-based_2017; oh_self-imitation_2018; imani_off-policy_2019; wilcox_monte_2022). In the offline context, geng_improving_2024 propose blending the MC return into the TD targets to achieve efficiency and stability benefits. While their motivation is similar, since they do not have a separate pre-training phase, the MC values associated with the behavior policy will become outdated as the policy improves, which can limit final policy performance. $Q$-Transformer (chebotar_q-transformer_2023) similarly claimed improved training efficiency at scale by incorporating MC values into the Bellman target via a max operation (as proposed by wilcox_monte_2022). The motivation is again comparable to ours, but since there is no separate pre-training phase, performance may be more affected by variance in the MC values, as we find in Appendix F. Additionally, this modification is only a secondary contribution of $Q$-Transformer that is not extensively investigated in isolation. We aim to more comprehensively understand and justify the use of MC targets for improving the efficiency of offline RL in this work.
3 Motivational Example
@@
Figure 1: A motivational tabular MDP. In offline reinforcement learning, we are provided with a dataset of trajectories. In this paper we utilize information from the entire trajectory (in the form of MC returns) to initialize a critic for subsequent off-policy reinforcement learning, which eliminates much of the inefficiency and instability associated with bootstrapping in temporal difference losses.
@@
As a motivational example, we consider offline tabular $Q$-learning on the simple 4 state MDP illustrated in Figure 1, together with a single offline trajectory from an unknown policy. In this example, we initialize all $Q$-values to zero (equivalent to randomly initializing a $Q$-network such that all initial $Q$-values are close to zero as is typical for deep $Q$-learning), and then perform temporal difference updates using the following update rule (sutton_reinforcement_2018):
@@
0 0 0 0 0 0 or +0.5111The MC initialization value for $Q(0,\rightarrow)$ depends on the choice of first-visit or every-visit Monte-Carlo (sutton_reinforcement_2018). In this example it doesn’t matter which is used after initialization. 0 +1 +3
@@
This minimal example demonstrates the inefficiency of using uninformed initial values even in the absence of function approximation, near-optimal behavior policies, or complex MDPs. Equivalent inefficiency in bootstrapping from randomly initialized temporal difference (TD) targets will occur in the case of uninformed neural network initializations, particularly in sparse-reward environments where rewards received at the end of long trajectories may take many TD updates to propagate.
4 Theoretical Analysis
@@
In this section we provide a more formal analysis of the intuition provided by the minimal tabular $Q$-learning example above. We aim to investigate the functional form with which the $Q$-function initialization can affect the convergence to optimal $Q$-values. We consider the case of approximate fitted $Q$-iteration (Riedmiller2005NeuralFQ), in which:
[…]
@@
where we have defined $\hat{Q}_{k}$ as the estimate of the $Q(s,a)$ function after $k$ iterations of Equation 3, and $\hat{T}$ as the approximate Bellman operator $\hat{T}Q=\hat{r}+\gamma\hat{P}\max_{a}Q$, where $\hat{r}(s,a)$ and $\hat{P}(s^{\prime}|s,a)$ are the sampled reward and environment transition probabilities respectively. The exact equivalents are denoted without the circumflex. We consider minimizing the infinite norm for each iteration since there is no convergence guarantee for the squared norm commonly used for empirical loss functions (agarwal_reinforcement_2022). In general, the error due to each $Q$-iteration (between successive iterations of the $\hat{Q}$ function) can be bounded as the sum of an ‘approximation’ error, and a ‘sampling’ error:
[…]
@@
This provides a measure of the (inifite-norm) Bellman error after $k$ iterations of fitted $Q$-iteration. We assume that for final convergence to the optimal policy, we would like to achieve an error over the offline dataset of order $\delta$. Therefore for two different intializations of $Q$, $Q_{0}$ and $Q^{\prime}_{0}$, we assume that it takes $k$ and $k^{\prime}$ iterations respectively to reach this error level, and make the bound tight:
[…]
@@
[…]
We see that $\hat{Q}^{\prime}_{0}\rightarrow\hat{Q}_{0}\implies k^{\prime}\rightarrow k$, i.e. a given initialization requires a given number of fitted $Q$-iterations to converge, as expected. Notice also the dependency on $1/\ln\gamma$, supporting our intuition that the efficiency improvement will generally be greatest for long-horizon tasks. We also see that as $\hat{Q}^{\prime}_{0}$ becomes closer to $Q^{*}$ (in terms of the infinite norm over all states and actions), the numerator grows with respect to the denominator and so $k-k^{\prime}$ increases logarithmically, demonstrating that the number of fitted $Q$-iterations $k^{\prime}$ decreases as the $Q^{\prime}_{0}$ initialization improves.
[…]